OpenStudy (anonymous):

how do i solve 5x+6-3x+8

OpenStudy (anonymous):

solve for x :) get x alone and the values alone : 2x = -8-6 x = -14/2 x = -7 ^_^

OpenStudy (anonymous):

is the equation equal to something? otherwise you cant solve it.

OpenStudy (anonymous):

yes i think it is

OpenStudy (anonymous):

you can add & subtract like numbers. So take the 5x and subtract the 3x (because it has a minus sign before it. Then add the 6 and 8. The result is your answer--you should have 2x+14

OpenStudy (anonymous):

you have just provided the left hand side of the equation. What is it equal to?

OpenStudy (anonymous):

is it asking to solve for x? the answer doesn't always have to be x=...unless it says to solve for x

OpenStudy (anonymous):

then it would have said simplify the expression, when you say solve, it means solve the equation. otherwise there is no meaning to it.

OpenStudy (anonymous):

Then set the entire problem to zero--bring the 6 & 8 over to the right side by subtracting. then divide out the x side

OpenStudy (anonymous):

Then set the entire problem to zero--bring the 6 & 8 over to the right side by subtracting. then divide out the x side